How much does a 0.40m long copper rod with a 6.7mm diameter stretch when one end is fixed and the other is pulled with a 1.1kN force?
Answer #1


Youngs modlulus Y of a material is given Y = FL/ADL

A = area

Y of copper = 1.1 e 11

DL = change of length

L = original length

F = force

so here we apply

DL = FL/AY

DL = 1100 * 0.4/(1.1e11 * 3.14* (6.7/2)^2 *10^-6)

DL = 0.113 mm
Answer #2

youngs modulus

?L=(FL)/(AY)

?L=[(1100N)(0.4 m )]/[(pi(.00335)^2)(117x10^9 )

?L=1.046x10^-4m